OpenStudy (anonymous):

a farmer sold 30 oranges and 20 apples for 50 dollars. if the apples cost 25 dollars. then what was the cost of each orange? show me the algebraic equation.

OpenStudy (anonymous):

25/30 = 5/6 dollar

OpenStudy (anonymous):

30x + 25 = 50 30x = 25 x = 25/30 = 5/6

OpenStudy (anonymous):

5/6 of a dollar is 83 cents 5/6(100) = 83